OpenStudy (anonymous):

The total cost (in thousands of dollars) of a machine manufacturing unit is given by the function T(x) = 300 + 5x, where x is the number of machines produced. What is the average cost (in thousands of dollars) of manufacturing 200 machines? (Average cost is the ratio of total cost to the number of machines manufactured.) 7.5 6.5 5.5 4.5 8

OpenStudy (perl):

average cost = T(x) / x

OpenStudy (perl):

lets call it A(x) A(x) = T(x)/x A(200) = T(200) / 200

OpenStudy (perl):

and there's no response...